What if initially a charge 2.5 C is put on shell #1 with radius  4.7 m, then a far away shell #2 (initially neutral) with radius  7.8 m is connected to shell #1 by a long conducting wire.

(a) What is the final charge on shell #1?
Q =  C


(b) What is the final charge on shell #2?
Q =  C


(c) What is the electric potential V on shell #1?
V =  V


(d) What is the electric potential V on shell #2?
V =  V
Answer #1

charge will flow until the potentials become same

a)

V1= v2

k*q1/r1 = k*q2/r2


q1/r1 = (Q-q1)/r2

q1/4.7 = (2.5-q1)/7.8

q1 = 0.94 C


b) q2 = Q-q1 = 1.56 C


c) V1 = k*q1/r = 1.8*10^9 V


d) V1 = v2 = 1.88*10^9 V